# A Network Science Approach to Granular Time Series Segmentation

### v1

@@ 5.1 Proposed GNN model
(b) Baseline sequence2point network architecture
@@
To realize the function $F_{G}()$, from Eq.3 to predict target classes from the transformed graph data denoted as $G$, a Graph Neural Network (GNN) was designed based on the Graph Attention Networks (GAT)[29]. The main advantage of GAT is a self-attention mechanism on node features that significantly improves classification performance for node classification tasks. This allows the model to calculate attention scores between nodes and their neighbors, which determine the influence each neighboring node has when updating a node’s feature representation. Unlike traditional GNNs, which aggregate information from all neighbors equally, GAT learns to focus on the most relevant ones, improving the accuracy of its predictions.
@@
As shown in Figure  2(a), the proposed architecture consists of 4 GAT layers. The first 3 layers have 32 neurons each but vary in the number of self-attention heads. The first two layers employ 4 attention heads, while the third layer employs 8 heads. The fourth layer has as many neurons as there are target classes and operates with 6 self-attention heads. We designed the architecture empirically by searching for the best ratio between the models complexity and performance. The configuration of the layers, filters, and attention heads was optimized through a grid search process, aiming to achieve the optimal ratio between model complexity and performance. The output of the final layer is passed to the output layer $Y$. For each of the nodes the network returns the probability of it being in one of the TS segments. All hidden layers use the Exponential Linear Unit (ELU) function for activation, while the finasl layer uses a softmax function. The ELU activation function handles the negative input values from the TS, while also contributing to the model convergence and accuracy.
5.2 Selected baseline model
@@
SotA DL TSS approaches, such as seq2seq [8], focus on sliding window methods, limiting granularity due to their fixed size. To overcome this and allow fair comparison to our proposed dense labeling approach, our baseline is based on the sequence2point (seq2point) DL model [30], which is an adaptation of seq2seq. This version avoids window-based segmentation constraints by predicting a single midpoint element of a window.
@@
The architecture of the baseline DL model is depicted in Figure 2(b). The input TS window is fed into five consecutive one dimensional convolutional layers. First layer consists of 30 kernels of size 10, second of 30 kernels of size 8, third of 40 kernels of size 6, followed by two layers consisting of 50 kernels of size 5. The output of the last convolutional layer is fed into a Fully connected or Dense layer consisting of 1024 node units, which is then connected to the output layer of size $Y$, where the softmax activation function is applied and the classification is made. Size of the $Y$ depends on the number of different classes present in the used dataset. All convolutional layers have a stride of 1, while both convolutional and Dense layers employ the Rectified Linear Units (ReLU) activation function.
@@ 6.1 Dataset
We validated our method using the TSSB dataset [14]. This dataset contains 75 annotated univariate TS with between 1 and 9 segments. Following common evaluation methodologies in TSS, we excluded TS containing only one segment, as they inherently lack change points and thus do not provide meaningful information about our method’s segmentation performance. Furthermore, TS containing duplicated segments were also discarded, since the dataset only contains the change point locations without the labels of each segment. As a result, our final dataset included 59 different TSSB datasets covering various domains, ensuring that our method is domain agnostic.
@@ 6.2 Model training and evaluation
Given that our datasets are imbalanced, where the number of instances in different classes is not equal, we assigned weights to the classes during the training process. The weights were calculated based on the inverse proportion of each class in the dataset. By assigning higher weights to the minority classes and lower weights to the majority classes, the model is forced to pay more attention to the minority ones. Otherwise it gets higher penalties if miss-classifying them. The models were trained for 1500 epochs using batch size 1 to process the graph as a whole.
[…]
@@ 7 Results
In this section, we evaluate the performance of our proposed TSS approach. The details of the experimental setup used to obtain the results are described in Section 6.
We start with a comparison of the baseline model with our proposed GAT model combined with the TS to graph transformations previously listed in Table 1. The utilized TSS methods were evaluated on the full set of 59 TSSB datasets, with the results visualized by boxplots in Figure 3. Each boxplot represents a specific combination of the proposed GAT model and a transformation. Different transformations are distinguished by color, with the color legend provided at the bottom of Figure 3.
Afterwards, we analyze how the models perform on datasets grouped by the number of segments. Figures 4-8 show the results for datasets containing between 2 and 7 segments and follow the same boxplot format as in Figure 3
At the end of the section, we explain the impact of different test-train split ratios on both the proposed and the baseline model.
@@ 7.1 Overall model performances
Figure 3 shows the overall performance of all implemented TSS methods across 59 univariate TSSB datasets. Each boxplot summarizes the distribution of performance scores for a given TS to graph transformation combined with the proposed GAT model, except for the first boxplot, which represents the baseline model. From left to right, the next 22 boxplots correspond to different VG transformations, followed by 5 boxplots for Transition networks and the last one for a Proximity network. The central line in each boxplot is the median, while the top and bottom edges mark the 75th and 25th percentiles, forming the interquartile range (IQR). The whiskers extend up to 1.5 times the IQR, and all points outside the whiskers are considered as outliers. In this context, higher medians and narrower boxes generally indicate stronger and more consistent performance. Of all the models evaluated, the WDPVG, the simple NVG and the weighted NVG variant with different edge weight initializations achieved the best results.
@@
If we look at the first boxplot from the left in Figure 3, we see a moderate performance with a median F1 score of 0.93. The IQR extends roughly from 0.9 to 0.97, which is slightly wider and suggests that the baseline model is less stable and more sensitive to fluctuations in the dataset compared to stronger NVG-based approaches. There are notable outliers at around 0.6 and 0.45, indicating lower performance on certain datasets.
@@
In contrast, the WDPVG method, second boxplot from the left in Figure 3, achieves the best overall performance compared to both the baseline and other TS to graph transformations. This is in line with our expectations outlined in Section 4, where we discussed the advantages of the WDPVG transformation, including its ability to capture additional structural information through its dual perspective. The median F1 score is positioned very high, close to 0.98. Combined with the narrow IQR and short whiskers, this suggests both superior performance and greater stability. This indicates that the WDPVG transformation consistently captures the underlying patterns and change points across various datasets, which underlines its suitability as the best performing transformation in combination with the GAT model.
@@
Comparing the undirected and unweighted NVG, shown in the third boxplot from the left in Figure 3 with the weighted NVG variants shown in boxplots 4-8, we observe a very similar overall performance. The median values across all variants are consistently above 0.9, with a similarly narrow IQR. This indicates that the topological or structural features have more influence on the segmentation process than the edge weights. Edges define the overall structure of the graph by connecting data points, while edge weights primarily influence how these relationships are interpreted rather than altering or influencing the topology itself.
@@
Boxplots 9-16 in Figure 3 represent the NVG variants with assigned edge directions. the left to right orientation outperforms the top to bottom direction, which is aligning naturally with the sequential order of TS data. Furthermore, we see that all directed variants exhibit higher variability and a greater number of outliers than their undirected counterparts. This is because directionality constrains the way the nodes in the graph are connected. In undirected VGs, every node is equally visible in both directions, whereas directed VGs filter out certain connections. As a result, the model misses out on useful structural information.
@@
Compared to the more successful NVG-based methods, the HVG variants listed in boxplots 17 to 23 from the left in Figure 3, exhibit a greater dispersion of performance scores and a higher number of outliers. This increased variability is not surprising, as the horizontal line of sight can oversimplify the relationships between data points, potentially overlooking important structural transitions. Consequently, the HVGs’ median performance values often fall below those of the NVGs, approximately around the value 0.85, and their wider IQRs and long extending whiskers suggest less consistency across different datasets.
@@
Finally, the MTF and proximity transformations, shown in boxplots 24 to 29 in Figure 3, resulted in a complex network representation that makes the detection of significant structures practically impossible in our TSS task. As illustrated in Figure 3, these transformations produce the lowest median values in their respective boxplots and have some of the widest IQRs compared to all other transformations. The probabilistic nature of these transformations not only captures the noise in the data, but also smooths out sudden changes or local fluctuations that are crucial for accurate segmentation [31].
@@
Figure 4: Boxplot summarizing the F1 score variability across various network architectures for 2 segment datasets.
@@
Figure 5: Boxplot summarizing the F1 score variability across various network architectures for 3 segment datasets.
@@ 7.2 Performance comparison based on number of segments
In this subsection, we compare the performance of our proposed WDPVG-based model with the baseline seq2point model across datasets with different number of segments. We specifically focus on the WDPVG, NVG, and baseline model, since NVG is foundational to WDPVG. Other models discussed previously have been analyzed in detail in Section 7.1.
Figure 4 shows the results for 21 TSSB datasets with two segments. In this simpler segmentation task, all three models perform similarly, as shown by the tightly clustered boxplots with median F1 scores just below 1.0. The corresponding Table 3 (in the Appendix section) shows that the baseline model achieves a slight lead overall, but the difference is minimal, and both WDPVG and NVG remain highly competitive in this task.
As the complexity increases to datasets with three segments, the WDPVG-based model starts to outperform the baseline model significantly, as shown in Figure 5. WDPVG achieves a median F1 score of around 0.98 with a narrow IQR of 0.97–0.99, indicating stable performance. In contrast, the baseline’s median drops to around 0.91, with a wider IQR (0.85–0.93), indicating less stability. NVG maintains its strong performance with the median value of approximately 0.95, placing it between WDPVG and baseline.
In Figure 6, the results for four segment datasets show the advantage of WDPVG even more clearly. The WDPVG model maintains its high performance with a median around 0.98 and an exceptionally low variability (IQR 0.97–0.985). The performance of the baseline model decreases, reaching a median value of around 0.9 with greater variability (IQR 0.87–0.94). NVG remains competitive, with a median performance close to 0.97, but slightly greater variability compared to WDPVG.
For datasets with five segments, WDPVG continues to outperform the competition, as shown in Figure 7. Again, the median F1 score is 0.98 and variability is minimal (IQR 0.965–0.985). NVG follows close behind and continues to perform strongly with a median score of 0.96, albeit with slightly greater variability. The baseline model, although slightly improved compared to the simpler scenarios, still lags behind (median approximately 0.94) and shows higher variability, including outliers around 0.85.
Finally, in the most complex scenarios with six or seven segments, illustrated in Figure 8, WDPVG shows exceptional performance, achieving a median of around 0.985 and consistently low variability (IQR 0.97–0.99). NVG also shows strong performance, though with slightly lower median values ( 0.97) and slightly higher variability. Conversely, the baseline model lags significantly behind both methods with a median of 0.91 and greater variability (IQR 0.87–0.93). These results suggest that the WDPVG is particularly effective for complex segmentation tasks, highlighting its ability to capture key structural dependencies within the data.

### v2

@@ 5.1 Proposed GNN model
(b) Baseline seq2point network architecture
@@
To realize the function $F_{G}()$, from Eq. 3 to predict target classes from the transformed graph data denoted as $G$, a Graph Neural Network was designed based on the Graph Attention Networks (GAT) [velivckovic2017graph]. The main advantage of GAT is its self-attention mechanism on node features, which significantly improves classification performance in node classification tasks. This mechanism enables the model to compute attention scores between nodes and their neighbors, determining the influence each neighboring node has when updating a node’s feature representation. Unlike traditional GNNs, which aggregate information from all neighbors equally, GAT learns to focus on the most relevant ones, improving the accuracy of its predictions.
@@
As shown in Figure 2(a), the proposed architecture consists of 5 GAT layers. The first four layers each have 32 neurons and four self-attention heads. The fifth layer contains as many neurons as there are target classes and operates with a single self-attention head. This specific five-layer architecture was empirically identified as the optimal configuration through our comprehensive Neural Architecture Search (NAS), described in Section 6.4. This search determined the best balance between model complexity and performance from a wide range of possibilities. The output of the final layer is passed to the output layer $Y$. For each node, the network returns the probability that it belongs to one of the TS segments. All hidden layers use the Exponential Linear Unit (ELU) activation function, while the final layer uses a softmax function. The ELU activation function manages the negative input values from the TS and contributes to model convergence and accuracy.
5.2 Selected baseline models
@@
SotA DL TSS approaches, such as seq2seq [phan_8631195], focus on sliding window methods, limiting granularity due to their fixed size. To overcome this and allow fair comparison to our proposed dense labeling approach, our baseline is based on the sequence2point (seq2point) DL model [zhang2018sequence], which is an adaptation of seq2seq. This version avoids window-based segmentation constraints by predicting a single midpoint element of a window.
@@
The architecture of the baseline DL model is depicted in Figure 2(b). The input TS window is fed into five consecutive one-dimensional convolutional layers. First layer consists of 30 kernels of size 10, second of 30 kernels of size 8, third of 40 kernels of size 6, followed by two layers consisting of 50 kernels of size 5. The output of the last convolutional layer is fed into a fully connected or Dense layer consisting of 1024 node units, which is then connected to the output layer of size $Y$, where the softmax activation function is applied and the classification is made. The size of the $Y$ depends on the number of different classes present in the used dataset. All convolutional layers have a stride of 1, while both convolutional and Dense layers employ the Rectified Linear Units (ReLU) activation function.
To benchmark our approach against SoTA unsupervised techniques, we employ the CLaP (Time Series State Detection) method [ermshaus2025clap]. Unlike simple CPD, which only localizes structural breaks, CLaP addresses the broader task of Time Series State Detection (TSSD): identifying and localizing recurring latent states within unannotated time series.In our implementation, we realize this state detection pipeline in two stages. First, we utilize ClaSP [ermshaus2023clasp], a parameter-free change point detector that exploits self-supervised learning to identify optimal segment boundaries based on classification accuracy. Once the time series is partitioned into distinct segments, we perform the state identification step by extracting a vector of statistical features for each segment (mean, standard deviation, minimum, maximum, and linear slope). These segment descriptors are then grouped using K-Means clustering to assign a discrete semantic state label to each segment. To ensure a rigorous evaluation of the method’s descriptive capability, the number of clusters $k$ is set to match the ground truth number of classes for each dataset.
@@ 6.1 Dataset
We validated our method using the TSSB dataset [ermshaus2023clasp]. This dataset contains 75 annotated univariate TS with between 1 and 9 segments. Following common evaluation methodologies in TSS, we excluded TS containing only one segment, as they inherently lack change points and thus do not provide meaningful information about our method’s segmentation performance. Furthermore, TS containing duplicated segments were also discarded, since the dataset only contains the change point locations without the labels of each segment. As a result, our final dataset included 59 different TSSB datasets covering various domains, ensuring that our method is domain agnostic.
@@ 6.2 Model training and evaluation
Given that our datasets are imbalanced, where the number of instances in different classes is not equal, we assigned weights to the classes during the training process. The weights were calculated based on the inverse proportion of each class in the dataset. By assigning higher weights to the minority classes and lower weights to the majority classes, the model is forced to pay more attention to the minority ones. Otherwise, it gets higher penalties if it misclassifies them. The models were trained for 1500 epochs using batch size 1 to process the graph as a whole.
For evaluating performance, we relied on the standard F1 score for each class, calculated as: $\textrm{F1 score}=\frac{2\times\textrm{Precision}\times\textrm{Recall}}{\textrm{Precision}+\textrm{Recall}}$, where: $\textrm{Precision}=\frac{\textrm{TP}}{\textrm{TP}+\textrm{FP}}$, $\textrm{Recall}=\frac{\textrm{TP}}{\textrm{TP}+\textrm{FN}}$, and TP, FP and FN stand for true positives, false positives and false negatives.
@@ 6.3 Comparative Analysis of Training Sample Ratios
6.4 Ablation study
To understand the nuanced contributions and significance of individual architectural components, we conducted a rigorous ablation study. Adopting a methodical approach, this study was structured as a comprehensive Neural Architecture Search (NAS) to systematically discern the relative influence of key hyperparameters on the trade-off between model performance and computational complexity. The NAS was implemented as a multi-objective optimization using the Optuna framework’s NSGAIISampler, simultaneously aiming to minimize the total number of model parameters and maximize the Mean F1-Score. This process enables the identification of a Pareto front of optimal solutions, revealing the functional dependencies between model size and classification accuracy.
The search space was designed to thoroughly investigate the model’s structural limitations. We evaluated network depth by varying the number of GATConv layers from 2 to 6, and network width by sampling hidden dimensions from 32, 64, 96, 128, 160, 192. To determine the optimal node aggregation strategy, we incorporated a comparison of attentive and non-attentive mechanisms directly into the search space. We examined the complexity of the attention mechanism by varying the number of GAT heads from 2, 4, 6, 8, while also evaluating a structural baseline in which the GAT layers were replaced with standard Graph Convolutional Network (GCN) layers. By including both architectures in the same optimization process, we could directly benchmark the value of dynamic neighbor weighting against static convolution under identical training constraints.
Additionally, the NAS optimized key training hyperparameters, including the learning rate, sampled from a log-uniform distribution between $5\times 10^{-4}$ and $3\times 10^{-3}$, and the weight decay regularization, chosen from a categorical selection of $\{0.0,1\times 10^{-4},1\times 10^{-3}\}$. All configurations were trained and evaluated on our comprehensive global panel, which consists of 59 time series.
@@ 7 Results
In this section, we evaluate graph-based frameworks for TSS, using the experimental setup detailed in Section 6. We first assess the overall viability of the approach by benchmarking the GAT model, combined with the transformations listed in Table 2, against two baselines: the unsupervised ClaSP + K-Means and the supervised seq2point model. These aggregated results across the full 59-dataset TSSB benchmark are presented in Figure 3.
Second, we examine how model performance correlates with the intrinsic complexity of the time series. To test this, we analyze the results by grouping datasets according to their number of segments, from 2 to 7, as shown in Figures 4-8. Finally, we assess generalization under data-constrained conditions by analyzing the impact of varying train-test split ratios on the leading WDPVG transformation relative to both baselines.
@@ 7.1 Overall model performances
Figure 3 shows the overall performance of all implemented TSS methods across 59 univariate TSSB datasets. Each boxplot summarizes the distribution of performance scores for a given TS to graph transformation combined with the GAT model, except for the first two boxplots, which represent the baseline models. From left to right, the next 22 boxplots correspond to different VG transformations, followed by 5 boxplots for Transition networks, and the last one for a Proximity network. The central line in each boxplot represents the median, while the top and bottom edges indicate the 75th and 25th percentiles, forming the interquartile range (IQR). The whiskers extend up to 1.5 times the IQR, and all points outside the whiskers are considered outliers. In this context, higher medians and narrower boxes generally indicate stronger and more consistent performance. Of all the models evaluated, the WDPVG, the simple NVG, and the weighted NVG variants achieved the best results.
The first boxplot from the left in Figure 3 represents the unsupervised ClaSP + K-Means baseline. While it achieves a competitive median score, the distribution reveals significant instability. The box is considerably wider than the top-performing graph methods, and the bottom whisker extends deeply, indicating large failures on specific datasets. This performance dispersion highlights the limitation of statistical baselines. While they excel at detecting clear trend changes or amplitude shifts, they struggle to generalize across datasets characterized by complex texture or subtle morphological changes.
@@
The second boxplot represents the seq2point baseline. It shows a relatively good performance with a median F1 score of approximately 0.93. The IQR extends roughly between 0.85 and 0.98. Compared to the stronger NVG-based approaches, this baseline model is less stable and more sensitive to fluctuations in specific datasets, as evidenced by notable outliers around 0.8 and 0.55.
@@
In contrast, the WDPVG method, shown in the third boxplot, achieves the strongest overall performance compared to both baselines and other graph transformations. This aligns with the expectations outlined in Section 4, regarding the advantages of the WDPVG transformation, specifically its ability to capture additional structural information through its dual perspective. The median F1 score is positioned very high, close to 0.98. Combined with the narrowest IQR and short whiskers, this suggests both superior performance and exceptional stability. This indicates that the WDPVG transformation consistently captures underlying patterns and change points across diverse signal domains, marking it as the most robust transformation for the TSS task.
@@
Comparing the undirected and unweighted NVG, shown in the fourth boxplot, with the weighted NVG variants shown in boxplots 5-8, we observe a very similar overall performance. The median values across all variants are consistently above 0.9, with a similarly narrow IQR. This suggests that topological or structural features have more influence on the segmentation process than edge weights. Edges define the overall structure of the graph by connecting data points, while edge weights primarily influence how these relationships are interpreted rather than altering the topology itself.
@@
Boxplots 9-16 in Figure 3 represent the NVG variants with assigned edge directions. The left-to-right orientation outperforms the top-to-bottom direction, aligning naturally with the sequential order of TS data. Furthermore, we see that all directed variants exhibit higher variability and a greater number of outliers than their undirected counterparts. This is because directionality constrains the way the nodes in the graph are connected. In undirected VGs, every node is equally visible in both directions, whereas directed VGs filter out certain connections. As a result, the model misses out on useful structural information provided by the backward visibility.
@@
Compared to the more successful NVG-based methods, the HVG variants listed in boxplots 17 to 23 exhibit a greater dispersion of performance scores and a higher number of outliers. This increased variability is not surprising, as the horizontal line of sight can oversimplify the relationships between data points, potentially overlooking important structural transitions. Consequently, the HVGs’ median performance values often fall below those of the NVGs, approximately around the value 0.85, and their wider IQRs and long extending whiskers suggest less consistency across different datasets.
@@
Finally, the MTF and proximity transformations, shown in boxplots 24 to 29, resulted in a complex network representation that makes the detection of significant structures practically impossible in our TSS task. As illustrated in Figure 3, these transformations produce the lowest median values and have some of the widest IQRs compared to all other transformations. The probabilistic nature of these transformations not only captures the noise in the data but also smooths out sudden changes or local fluctuations that are crucial for accurate segmentation [wang2015encoding].
@@
Figure 4: Boxplot summarizing the F1 score variability across various network architectures for 2-segment datasets.
@@
Figure 5: Boxplot summarizing the F1 score variability across various network architectures for 3-segment datasets.
@@ 7.2 Performance comparison based on number of segments
In this subsection, we evaluate the performance of the leading graph-based transformation WDPVG against both the supervised seq2point baseline and the unsupervised ClaSP + K-Means baseline across datasets with varying numbers of segments. We specifically focus on the WDPVG, NVG, and baseline model, since NVG is foundational to WDPVG. Other models discussed previously have been analyzed in detail in Section 7.1.
Figure 4 illustrates the results for the 21 datasets containing two segments. In this simpler binary segmentation task, the performance landscape is mixed. The ClaSP + K-Means baseline, shown in the first boxplot, demonstrates high potential with a median near 1.0 but suffers from extreme instability, indicated by a very wide IQR and a whisker extending down to 0.5. This suggests it either solves the task perfectly or fails significantly depending on the dataset. The seq2point baseline shows more stability but a slightly lower median. Both WDPVG and NVG, represented as the third and fourth boxplots, exhibit highly competitive and stable performance, with medians approaching 1.0 and very narrow IQRs, proving their robustness even in simple scenarios.
As complexity increases in datasets with three segments, the graph-based methods begin to distinguish themselves, as shown in Figure 5. WDPVG achieves a dominant median F1 score of approximately 0.98 with a tight IQR (0.95–0.99), indicating superior stability. The ClaSP baseline shows a low median, and remains the most volatile method, with a boxplot that spans a wide range of performance scores. The seq2point baseline drops, with a median around 0.92 and increased variability. NVG maintains strong performance with a median near 0.98, positioning it as a solid intermediate between the baselines and WDPVG.
n Figure 6, the results for four-segment datasets further highlight the robustness of WDPVG. The WDPVG model maintains exceptional performance with a median of roughly 0.99 and negligible variability (IQR 0.98–1.0). The ClaSP baseline shows improved stability compared to the two- and three-segment scenarios. iT achieves a high median near 1.0 and a tighter interquartile range, though it still exhibits a long lower whisker extending down to approximately 0.68, indicating occasional underperformance. The seq2point baseline performs reliably with a median around 0.95, but displays a wider spread than the graph-based methods. NVG remains highly competitive, mirroring WDPVG’s high median performance, albeit with a slightly increased number of outliers.
For datasets with five segments (Figure 7), performance is highly competitive across all methods. WDPVG and NVG remain top performers, both achieving medians effectively at 1.0 with negligible variance. Notably, the ClaSP + K-Means baseline also demonstrates exceptional stability in this specific category, matching the graph methods with a median near 1.0 and a narrow IQR, though it exhibits outliers down to 0.8. The seq2point baseline shows marked improvement, reaching a median close to 0.99 with a tight distribution, yet similarly suffers from distinct performance drops.
Finally, in the most complex scenarios with six or seven segments (Figure 8), WDPVG demonstrates its scalability, achieving a median of nearly 0.99 with consistently low variability (IQR 0.98–1.0). NVG also performs strongly, though with slightly lower consistency than WDPVG. Conversely, the seq2point baseline falls behind with a median of 0.88 and significant variability. The ClaSP baseline exhibits its characteristic instability, with a high median but a distribution that suggests frequent failures. These results confirm that WDPVG is particularly effective for complex, multi-segment tasks, capturing structural dependencies that purely statistical or sequential baselines often miss.